A sphere A with charge of +6 micro culombs is located near B of-3micro culombs is located 4cm to the right of A... What is the force of sphere B on A? 

Grade:12th pass

1 Answers

Khimraj
5 years ago
F = KQq/r2 = 9*109*6*10-6*(-3*10-6)/16*10-4 = -101.25 N
Hope it clears.....................................................................